OpenStudy (anonymous):

It takes a man 8 hours to wash and wax the family car. If his teenage son helps him, it only takes 5 hour(s). How long would it take the son, working alone, to wash and wax the car? @Calculus1

OpenStudy (anonymous):

8 (rate of the man) = 5 (rate of the man + rate of his son) = 1 rate of the man = 1/8 rate of his son = 3/40 3/40 (time) = 1 time = 40/3 hour

OpenStudy (lilg132):

where did u get 40 from?